# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v4

@@ 4.2 ILP formulation
These are introduced to bind the train with the shunting movement called also service train. If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
$\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s).$ (14)
@@
We have additional conditions on the $y$-variables, concerning the case when M-O is not possible on the line or station.
@@
$\begin{split}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}\\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\\ y^{(out)}(j,j^{\prime},s)=y^{(out)}(j,j^{\prime},s^{\prime})\\ \forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\\ y^{(in)}(j,j^{\prime},s^{\prime})=y^{(out)}(j,j^{\prime},s^{\prime})\end{split}$ (15)
The objective function in Eq. (2), together with the constraints in Eq. (3)-(14) define our mathematical programming model.
The model yields an integer linear program, and the time variables can be constrained even into a finite range using the parameter $d_{\max}$. Note that the binary variables have the obvious symmetry property
@@
$\begin{split}y^{(\text{out})}(j^{\prime},j,s)=1-y^{(\text{out})}(j,j^{\prime},s)\\ z(j^{\prime},j,s^{\prime},s)=1-z(j,j^{\prime},s,s^{\prime}),\end{split}$ (16)
which is taken into account explicitly by using independent binary variables only.
@@
As for the scaling of our model, the number of time variables i.e. #$\#t$ is bounded by number of trains times number of stations #$\#\mathcal{J}\#\mathcal{S}$ (see also Section 3.1 of the work of  Domino et al. (2022b)). To obtain a more accurate estimate of the number of these variables we observe that in our model no trains visit all stations. We assume that in average each train visits $\alpha\in(0,1)$ fraction of the stations, we have then
#$\#(t)\approx\alpha\#\mathcal{J}\#\mathcal{S}.$ (17)
as a tighter bound.
Suppose, that most of the network is composed mainly of double-track lines. Then there will be typically a single precedence variable per train and station ($y^{(out)}(j,j^{\prime},s)$), and thus their number can be estimated by:
#$\#y+\#z\approx\#y\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S},$ (18)
where $N(d_{\max})$ is the number of trains the train can meet at a station on average. Under such assumptions, the total number of variables grows linearly with number of trains and stations:
#$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}\#\mathcal{S}.$ (19)
If, on the other hand, the network has dominantly single-track lines, two precedence variables per train and station ($y^{(out)}(j,j^{\prime},s)$ and $z(j,j^{\prime},s,s^{\prime})$) will be needed, and thus
#$\#y=\#z\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S}.$ (20)
In both cases we assume that there are not many ($y^{(in)}(j,j^{\prime},s)$) variables, as they are tied certain particular interlocking area conditions.
Based on Eq. (16) there are typically $2$ constraints per each $y$ variable as headway and station track occupation constraints, and $2$ constraints per each $z$ variable as single-track occupation constraints. For the interlocking area occupation constraints, it is more complicated as it involves both $z$ and $y$ variables; a good assumption is to consider $2$ constraints for each $y$ variable. For the running time and minimal dwell time we expect one constraint per train and station. This number of constraints can be estimated by
#$\#constr.\approx 6\#y+2\#z+2\#\mathcal{J}\#\mathcal{S}$ (21)
where the second term is tied to single-track line conditions (one condition per $z$ variable).
#$\#constr.\approx\begin{cases}\alpha(6N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for double-track}\\ \alpha(8N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for single-track}.\end{cases}$ (22)
Because of the limitations in time imposed by $d_{\max}$ the model can be viewed as linear in number of trains and number of stations. Hence, as expected, the model is more complex for the network in which single-track lines dominate. Although the analyzed network is mostly composed of the double-track lines, we will analyze certain use-cases with the fraction of single-track lines increased.
Observe that without limiting the secondary delays by $d_{\max}$, the value of $N(d_{\max})$ would only be bounded by #$\#\mathcal{J}-1$ (as each train is considered to possibly meet each other), which would result in quadratic scaling of the number of precedence variables, and that of the number of constraints in the number of trains.
@@ 5 Hybrid quantum-based approach
To overcome the limitations described in in Section 2 of current hardware quantum annealers, we apply a hybrid (quantum-classical) solver. In particular, we use the ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a); a cloud-based proprietary solver that is developed by the market leader of QA hardware and is available as a service. It utilizes a hybrid approach that combines classical computational power with quantum processing.
In particular, we have used the Constrained Quadratic Model (CQM) (D-Wave Quantum Inc., 2022b) Solver. The CQM hybrid solver works as follows.
@@
Preprocessing includes subproblem identification and decomposition to smaller instances (Tran et al., 2016). Furthermore, the solver’s internal preprocessing mechanism takes constraints into account automatically, implementing the suitable penalty method internally.
@@
The hybrid solver implements a workflow that combines a portfolio of various classical heuristics (including tabu search, simulated annealing, etc.) running on powerful CPUs (Central Processing Units) and GPUs (Graphics Processing Units). In the course of the solution process the hardware quantum annealer is also invoked, to solve smaller QUBO problems that can potentially boost the classical heuristics. Integer problems are transformed to binary using state-of-the-art methods like the one described by Glover et al. (2019); Karimi and Ronagh (2019), generally applicable for integer problems. When using the quantum hardware, the solver performs the required embedding, runs the QUBO subproblem on the physical QA device several times, and returns a sample of potential solutions. After these readouts, the sample is incorporated into the solution workflow. In this way, even though the physical annealer supports problems with limited size (e.g. the Pegasus machine has $5500$ physical qubits (Dattani et al., 2019), and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process. The final solution is composed by the solver’s module.
@@
The solution or solutions of quadratic constrained problem or the integer problem. The workflow results in a solution or solutions that are not guaranteed to be optimal, hence, the hybrid solver as a whole is in nature a heuristic solver. Furthermore, the solver is probabilistic, hence at each run the final output can be different. The CQM (integer) solver returns a series of solutions, which can be an additional asset. Nevertheless, for sake of presentation, our approach is to analyze only the best CQM solution given each run of the hybrid algorithm (realization).
@@
As mentioned before, in addition to the actual problem to be solved, the CQM solver optionally inputs another parameter, t_min. This is a time limit for heuristics running in parallel(D-Wave Quantum Inc., 2022b): unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution. We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting on the model performance and solution quality for our problem.
The CQM Solver can solve problems encoded in the form of Eq. (2) – (15) (c.f. Section 4.2) with up to $5000$ binary or integer variables and $100,000$ constraints. Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum-classical) solvers. As a classical solver, IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) was used. The CPLEX computation was performed on $16$ cores of Intel(R) Core(TM) i7-10700kF CPU 3.80GHz with $64$GBs of memory. As a hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used.
@@ 6 Computational results
In this section, we demonstrate the performance of the proposed hybrid (quantum-classical) approach on a network at Polish railways - the central part of the Upper Silesia Metropolis. In particular, we compare the performance of D-Wave’s hybrid quantum-classical solver (CQM) with CPLEX. The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different level of disturbances including various initial delays and track closures. As the objective is tied to secondary delays, we can roughly apply it to assess the degree of difficulty of the problem.
Section 6.1 describes the considered network and traffic characteristics. Section 6.2 presents results concerning generic examples on the core line of the network, comparing double-track-line scenarios with single-track-line scenarios and closure scenarios. Section 6.3 presents analogous results for larger network with real-life timetable. Here the case studies concern various scenarios with different degrees of difficulty.
@@ 6.1 Railway network
Figure 2: Selected part of Polish railway network located in the central part of the Metropolis and examined in our research.
The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analyzed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single-track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O). As described before, the decision stations are the only ones that appear explicitly in our model. Due to operational reasons resulting from track design, the decisions on train departure times (and thus on trains’ order) are, in the current operational practice, de facto made with respect to the decision stations. All other stations and junctions are treated as line blocks as long as no rerouting or retracking is considered; they will appear in the model implicitly via parameters. These stations include:
@@
In all computation priority weights in Eq. (2) we use $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable). As for the solver parameter t_min, after initial experiments, we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality. A sensitivity analysis regarding this parameter will be performed in Section 6.3.

### v5

@@ 4.2.2 Constraints
These are introduced to bind a train with a shunting movement, termed also as a service train. If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s),\ \forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}.$ (16)
@@
We have additional conditions on the $y$-variables, concerning the case when M-O is not possible on the line or station. In particular, assume that the trains $j$ and $j^{\prime}$ are following each other from $s$ to $s^{\prime}$, heading in the same direction. If they use the same track on $s^{\prime}$, they will have to leave $s^{\prime}$ in the same order as they left $s$. Formally:
@@
$\begin{split}y^{(\text{out})}(j,j^{\prime},s)=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(j_{0},j_{0}^{\prime})\in\mathcal{J}^{2\,(\text{headway})},\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j_{0},j_{0}^{\prime}}\\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\end{split}$ (17)
Similarly, if trains $j$ and $j^{\prime}$, enter $s^{\prime}$ from any location, they are passing the switching zone of station $s^{\prime}$, then they are passing the station $s^{\prime}$ in the same direction using the same track, they have to keep their order upon departure. Formally:
@@
$\begin{split}y^{(\text{in})}(j,j^{\prime},s^{\prime})=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(s,s^{\prime})\in S\times S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{\text{in}},\text{MO})}_{s^{\prime},s}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}.\end{split}$ (18)
The objective function in Equation (2), together with the constraints in Equation (3)-(16) define our mathematical programming model.
Our model is formulated as an integer linear program. Moreover, as the time variables are constrained a finite range using the parameter $d_{\max}$, all variables are of a finite range. We intend to limit the number of variables as much as possible. The binary variables have the obvious symmetry property
$\begin{split}y^{(\text{out})}(j^{\prime},j,s)=1-y^{(\text{out})}(j,j^{\prime},s)\\ z(j^{\prime},j,s^{\prime},s)=1-z(j,j^{\prime},s,s^{\prime}).\end{split}$ (19)
These are added to the model as constraints, in a way that only half of the variables are considered as independent, and the rest are removed.
4.2.3 Scaling analysis
@@
Let us discuss the number of variables and constraints in our model. The number of time variables #$\#t$, the product #$\#\mathcal{J}\#\mathcal{S}$ of the number of trains and the number of stations serves as a trivial but rough upper bound. (See also Section 3.1 of the work of  [17]). To obtain a more accurate estimate of the number of these variables we observe that in our model no train visits all the stations. To provide an estimate, we introduce $\alpha\in(0,1)$ as the fraction of the stations visited by an average train. Thus we have
#$\#(t)\approx\alpha\#\mathcal{J}\#\mathcal{S}\leq\#\mathcal{J}\#\mathcal{S}.$ (20)
as a tighter bound. In what follows we will mostly $\alpha=1$, the rough bound, however, we will keep the parameter $\alpha$ in the formulae to reflect its role in the scaling. Let us assume, that half the of trains (#$\frac{\#\mathcal{J}}{2}$) are going in one direction and the other half are going in the other. Then the number of trains each one meets going in the same direction is $\alpha\frac{\#\mathcal{S}\#\mathcal{J}}{2}$. This yields #$\#\mathcal{J}\cdot\frac{\alpha\#\mathcal{S}\#\mathcal{J}}{2}$ $y^{(\text{out})}$ variables. However, because of the symmetry property in Equation (19), only half of these variables are independent, thus we have $\alpha\#\mathcal{S}\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{out})}$ of these variables. As for the trains going in opposite directions, we assume that each train meets at most #$\frac{\#\mathcal{J}}{2}$ trains on each station in the single-track parts of the network. Hence, the number of $z$ variables is at most #$\#\mathcal{J}\cdot\frac{1}{2}\cdot\frac{\alpha\#\mathcal{J}(\#\mathcal{S}-1)}{2}$ (here, #$\#\mathcal{S}-1$ comes from the observation that the $z$-variables are tied to pairs of stations). Variables $y^{(\text{in})}$ occur mainly in Eqs. (4.2.2)-(4.2.2), relevant for a few stations only: there is typically one such station in the network, yielding at most #$\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{in})}$ variables. In summary, the number of precedence variables can be estimated as
#$\begin{split}&\#y^{(\text{out})}\leq\alpha\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\leq\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\\ &\#y^{(\text{in})}\leq\frac{\#\mathcal{J}^{2}}{4}\\ &\#z\leq\alpha\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\leq\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\\ &\#y+\#z=\#y^{(\text{out})}+\#y^{(\text{in})}+\#z<\frac{\#\mathcal{J}^{2}}{2}\#\mathcal{S}\end{split}$ (21)
Note that these bounds are rather pessimistic: setting the parameter $d_{\max}$ as an upper bound on the secondary delays implies a reduction in the number of trains that can potentially conflict. However, the actual number of them can depend on the train density and structure of the timetable in various ways, hence, it is not possible to provide a tighter bound that would be valid in general. Finally, let us remark as the $z$-variables are mainly tied to single-track traffic, the model is more complex for networks dominated by single-track segments.
Based on Eqs. (20) and Equation (21), the scaling of the upper bound on the the total number of decision variables can be estimated by
#$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}^{2}\#\mathcal{S}+\text{const'}\#\mathcal{J}\#\mathcal{S},$ (22)
which is quadratic in the number of trains and linear in number of stations.
Now let us estimate the number of constraints. Considering the headway constraints in Equation (4), there is one such constraint for each variable. However, because of the symmetry property in Equation (19), only half of the variables are independent, hence, the number of these constraints is twice the number of independent variables.
Along the same arguments, in case of the constraints in Equation (11), i.e. station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each $z$ variable. It is more complicated to estimate the number of interlocking area occupation constraints in Eqs. (4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables. A good assumption is to consider $2$ constraints for each $y^{(\text{out})}$ variable, $2$ constraints for each $y^{(\text{in})}$ variable, and $2$ constraints for each $z$ variable. Further, we expect one minimal dwell time constraint per train and station (described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term). Concerning train order constraints in Eqs (17) (18), we also expect one constraint per train and station (also described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term). We exclude the rolling stock circulation constraints from this consideration as their contribution to the number of constraints is very limited. In summary, the total number of constraints obeys the following upper bound:
#$\#constr.<6\#y^{(\text{out})}+2\#y^{(\text{in})}+4\#z+2\alpha\#\mathcal{J}\#\mathcal{S}.$ (23)
Substituting the estimated number of variables in Equation (21) we have
#$\#constr.<\alpha\frac{3}{2}\#\mathcal{J}^{2}\#\mathcal{S}+\frac{1}{2}\#\mathcal{J}^{2}+\alpha\#\mathcal{J}^{2}(\#\mathcal{S}-1)+2\alpha\#\mathcal{J}\#\mathcal{S}$ (24)
Setting $\alpha=1$ as the worst case, we get
#$\#constr.<\frac{5}{2}\#\mathcal{J}^{2}\#\mathcal{S}-\frac{1}{2}\#\mathcal{J}^{2}+2\#\mathcal{J}\#\mathcal{S}.$ (25)
Let us stress that the provided generic pessimistic upper bounds only serve the purpose of getting an overall picture of the scaling; the particular number of variables and constraints is lower and will be specified for our problem instances.
@@ 5 Hybrid quantum-based approach
To overcome the limitations of current hardware quantum annealers described in Section 2, we apply a hybrid (quantum-classical) solver. In particular, we use the ’Leap Hybrid Solver Service (HSS)’ [8]; a cloud-based proprietary solver that is developed by the market leader of QA hardware and is available as a service. It implements a hybrid approach that combines classical computational power with quantum processing. In particular, we have used the Constrained Quadratic Model (CQM) [9] Solver. The CQM hybrid solver works as follows.
@@
Preprocessing includes subproblem identification and decomposition to smaller instances [38]. Furthermore, the solver’s internal preprocessing mechanism takes constraints into account automatically, implementing the suitable penalty method internally.
@@
The hybrid solver implements a workflow that combines a portfolio of various classical heuristics (including tabu search, simulated annealing, etc.) running on powerful CPUs (Central Processing Units) and GPUs (Graphics Processing Units). In the course of the solution process the hardware quantum annealer is also invoked, to solve smaller QUBO problems that can potentially boost the classical heuristics. Integer problems are transformed to binary using state-of-the-art methods like the one described by [19, 24], generally applicable for integer problems. When using the quantum hardware, the solver performs the required embedding, runs the QUBO subproblem on the physical QA device several times, and returns a sample of potential solutions. After these readouts, the sample is incorporated into the solution workflow. In this way, even though the physical annealer supports problems with limited size (e.g. the Pegasus machine has $5500$ physical qubits [12], and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process. The final solution is assembled by the solver’s module.
@@
The solution or solutions of quadratic constrained problem or the integer problem. The workflow results in a solution or solutions that are not guaranteed to be optimal, hence, the hybrid solver as a whole is in nature a heuristic solver. Furthermore, the solver is probabilistic, hence at each run the final output can be different. The CQM (integer) solver returns a series of solutions, which can be an additional asset. For the probabilistic nature of the solver, the returned set of solutions is often termed as a sample, and the process of obtaining the sample is termed as sampling. Note that CQM, like most heuristic solvers, does not warrant feasibility by default, hence, a direct check is performed after obtaining the possible solutions. Nevertheless, for the sake of presentation, our approach is to analyze only the best CQM solution given each run of the hybrid algorithm (realization).
@@
As mentioned before, in addition to the actual problem to be solved, the CQM solver optionally inputs another parameter, t_min. This is a time limit for heuristics running in parallel[9]: unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution. We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting on the model performance and solution quality for our problem.
The CQM Solver can solve problems encoded in the form of Equation (2) – (19) (c.f. Section 4.2) with up to $5000$ binary or integer variables and $100,000$ constraints. Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum-classical) solvers.
As a classical solver, we have opted for IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) as one of the most prevalently used commercial solver. The CPLEX computation was performed on $16$ cores of Intel(R) Core(TM) i7-10700kF CPU 3.80GHz with $64$GBs of memory. As a hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used. In addition to solving the problems to optimality, we include CPLEX results with limited computational time, for the sake of a fair comparison with heuristics.
@@ 6 Computational results
In this section, we demonstrate the performance of the proposed hybrid (quantum-classical) approach on a network of Polish railways - the central part of the Upper Silesia Metropolis. In particular, we compare the performance of D-Wave’s hybrid quantum-classical solver (CQM) with CPLEX. The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different levels of disturbances including various initial delays and track closures. The considered closures are situations in which one or more tracks between decision stations are closed and the traffic has to be redirected to the remaining tracks between decision stations. This can imply the need to add extra decision stations to the model to bind the part of the line with the closure; this also occurs in our examples.
Section 6.1 describes the considered network and traffic characteristics. Section 6.2 presents results concerning generic synthetic examples on the core line of the network, comparing double-track-line scenarios with single-track-line scenarios and closure scenarios. Section 6.3 presents similar results for larger network with real-life timetable. Here the case studies concern various scenarios with different degrees of difficulty.
@@ 6.1 Railway network
Figure 2: The part of the Polish railway network which is the subject of our considerations. It is located in the central part of the Metropolis. The displayed non-decision stations such as ZZ or RCB become decision stations in certain problem instances, e.g. due to an assumed closure of one of the tracks between ZZ and RCB.
The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analyzed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single-track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O). As described before, the decision stations are the only ones that appear explicitly in our model by assigning decision variables. Due to operational reasons resulting from track design, the decisions on train departure times (and thus on trains’ order) are, in the current operational practice, de facto made concerning the decision stations. All other stations and junctions are treated as line blocks as long as no rerouting or retracking (e.g. in the case of closure) is considered; they will appear in the model implicitly via parameters. In such a case precedence of trains cannot be changed on these stations as there are no decision variables assigned to them.
These stations include:
@@
The priority weights in Equation (2) are set to $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable), in all the computations. These weights result from interpreting the instructions on train traffic management rules and priority levels for trains set out in the ”Instruction for train operation (Ir-1)” 222see the paragraph $51$ of Instrukcja o prowadzeniu ruchu pociagow Ir-1 (in Polish), Eng. Instruction for railway traffic https://www.plk-sa.pl/files/public/user_upload/pdf/Akty_prawne_i_przepisy/Instrukcje/Wydruk/Instrukcja_o_prowadzeniu_ruchu_pociagow_Ir-1_-_wchodzi_w_zycie_od_14_sty....pdf accessed 18 November 2024 and the real-life dispatching practices on the PKP PLK network. The default value for regular traffic is $w_{j}=1$, while there is no priority for shunting. Observe the higher priorities of InterCity trains and the even slightly higher priorities of express trains operated by Express Intercity Premium bullet trains which are briefly referred to as “express trains”. These intercity trains and express trains have strong priority in the case of small-scale disturbances, but this priority can be neglected in the case of larger disturbances. As for the solver parameter t_min, after initial experiments, we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality. A sensitivity analysis regarding this parameter will be performed in Section 6.3.